OpenStudy (lgbasallote):

another demonstration pls @Mimi_x3 and @apoorvk ? :DD \(\int \frac{dx}{x^2 \sqrt{4 - x^2}}\) i think it's an easy one but please show me steps :DDD Note: Trigonometric Substitution

OpenStudy (blockcolder):

To make the pesky radical disappear, let x=2sin(theta).

OpenStudy (lgbasallote):

uh-huh...then..?

OpenStudy (turingtest):

\[x=2\sin\theta\implies dx=2\cos\theta d\theta\]sub that in\[\int{2\cos\theta d\theta\over(2\sin\theta)^2\sqrt{4-4\sin^2\theta}}\]

OpenStudy (turingtest):

simplify...

OpenStudy (turingtest):

from the \(x=2\sin\theta\) sub we can infer this trianglewhich will help us determine our final expression